Operations Report – December 18th

Crew 185
Operations Report Monday 12/18/2017
SOL: -1
Name of person filing report: A.Passeron
Non-nominal systems:
– Heater :
On Monday morning, Thibault and I removed the air filter. It had not been changed since April 2016. At 10:00am, Ilaria and David went to city to buy a battery for the generator and a pack of air filters. We replaced the filter at 3:00pm. The heater is now working well.
-Generator :
Shannon helped us to change the battery at 3:00pm. We then drained oil from the radiator by pouring clean cooling liquid. The oil of the generator was also very dirty. It was so sticky that we had to start the generator for a few minutes to heat the oil and thin it. We thus were able to remove the oil filter, but only a fraction of dirty oil came out of the generator. We tried to remove as much dirty oil as possible by filling the motor with clean oil, but a valve seemed to block the oil circulation. After pouring as much clean oil as we could, we replaced the oil filter. The generator was finally turned on at 5:00pm.
Generator (hours run):
Generator turned on at 5:00pm.
Solar— SOC @ 6:00pm : 98%
Diesel – 80%
Propane – 60 percent volume
Ethanol Free Gasoline (5 Gallon containers for ATV) – 5 Gallons (Full)
Water (trailer) – 150 gallons
Water (static) – 300 gallons
Trailer to Static Pump used – Yes
Static to Loft Pump used – Yes
Water Meter: 45 gallons
Toilet tank emptied: No
ATV’s Used: No
Oil Added? No
ATV Fuel Used:
# Hours the ATVs were used today: 00:00 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Opportunity rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Curiosity rover used:  No
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: All Nominal
Summary of Hab operations: All nominal
Summary of GreenHab operations: All nominal
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support:

Sol Summary – December 18th

Today has been a busy day for the Crew 185. As it was our last day before entering simulation, we needed to prepare the hab in the best possible way and to fix as many things as possible before closing our doors.

We decided to work in two groups. Ilaria and David took the car to buy a new battery for the generator, soil for the greenhab and food. Arno, Thibault and John set the laboratory with the 3D printer, the medical supplies, the X-1 spacesuit prototype, etc…

This afternoon, we repaired the generator by replacing the battery and changing a part of the cooling liquid and the oil. It is working better at the moment when I write these lines! Hopefully, it will keep being in a good shape.

We decided that we would begin the simulation tomorrow morning when waking up. Mars, here we come!

The Crew 185

Copyright © The Mars Society. All rights reserved. | Main Site